In 2019, X Company sold 6,250 units of its only product. Total revenue was $1,608,750, total variable costs were $660,000, and total fixed costs were $642,000. In 2020, there are only two expected changes: variable costs will increase by $4.90 per unit, and fixed costs will decrease by $49,800. How many units must X Company sell in 2020 in order to earn $34,000?
Answer #1

in 2019

selling price per unit = 1608750 / 6250 = 257.4

variable cost pert unit = 660000 / 6250 = 105.6

fixed cost    = 642000

in 2020

selling price    = 257.4 per unit

   variable cost per unit = 105.6 +4.9 = 110.5

fixed cost    = 642000 - 49800 = 592200

contribution per unit = selling price per unit - variable cost per unit

   = 257.4 - 110.5

   = 146.9

required units to earn profit 34000 = ( fixed cost + profit ) / contribution per unit

= ( 592200+34000 ) / 146.9

   = 4262.76 units